How many microstates are accessible to a single water molecule at 298K?

Standard Molar Entropy of water is 188.8 J/(mol*K)
1
Expert's answer
2018-01-28T09:01:54-0500
The number of microstate, from which this macroscopic state of the system is composed, is called the thermodynamic probability W.

In practice, they deal with a huge number of particles in the system, which affects the large values of the thermodynamic probability; therefore, they do not use the absolute value of W, but its logarithm lnW.

S = R·lnW

S°298 = 188.8 J/(mol*K)

R = 8.314 J/(mol*K)

lnW = S/R = 188.8 J/(mol*K) / 8.314 J/(mol*K) = 22.7

6.02*1023 molecules - 22.7

1 molecule - lnW

lnW = 22.7/6.02*1023 = 3.77*10-23